OpenStudy (anonymous):

can this simplify anymore? +-2√20 over 12

OpenStudy (anonymous):

sqrt(5)/3

OpenStudy (anonymous):

\[\frac{\pm \sqrt{5}}{3}\]

OpenStudy (anonymous):

20 is a multiple of 4 and 5. 4 is a perfect square. Therefore, \[{2*\sqrt{4*5} \over 12} = {4*\sqrt{5} \over 12} = {\sqrt{5} \over 3}\]

OpenStudy (anonymous):

Thank You!!

jhonyy9 (jhonyy9):

eashmore where is +/- ?